OpenStudy (anonymous):

If w(t) = 3t + 5, solve w(t - 1) = w(2t).

OpenStudy (loser66):

w(t) = 3 t +5 w(t-1) = 3(t-1) +5 = 3t+2 w(2t) = 3(2t) +5 w(t-1) = w(2t) you just let them equal and solve for t